During exercise, both QO2 and QCO2 increases as well as cardiac output, Qa. Suppose that during exercise QO2 increases to 2000 mL/min and the respiratory quotient remains at 0.8, and that Qa increases to 18 L/min. Assume that PaO2 remains at 95 mmHg, that O2 saturation is at 98% and PaO2 is 40 mmHg and that blood [Hb] = 15 g %.

A. What is the total arterial content of O2?

B. What is the venous content of O2?

C. What is the PVO2?

D. What is Qco2?

E. Assuming an arterial pH is 7.4 what is the total CO2 cocnt of arterial blood?

F, Can you determine the toal CO2 content of venous blood?

G. Can you determine PVCO2?

H. Determine the new alvelor ventilation from alvelor ventilation equation

I. Calculate the predicted PaO2 from alvelar gas equation
Answer #1

Given RQ=0.8, QO2= 2000ml/min

QCO2= RQ*QO2

= 0.8* 2000

=1600ml/ min

1) Given Hb=15% , SaO2=98, PaO2=95

CaO2=(1.34*Hb*SaO2) + (0.0031*PaO2)

=(1.34*15*98)+(0.0031*95)

=1970.0945ml

2) Given Hb=15, SvO2=98, PvO2=40

CvO2=(Hb*1.36)(SvO2/100) +(PvO2*0.0031)

=(15*1.36)(98/100) +(0.0031*40)

=19.992+0.124

=20.116ml

3). Given that the PvO2 is 40mm/Hg

4). QCO2= RQ*QO2

=0.8*2000

=1600ml/min